# Scene Flow as a Partial Differential Equation

### v1

@@ Abstract
We reframe scene flow as the task of estimating a continuous space-time PDE that describes motion for an entire observation sequence, represented with a neural prior. Our method, EulerFlow, optimizes this neural prior estimate against several multi-observation reconstruction objectives, enabling high quality scene flow estimation via pure self-supervision on real-world data. EulerFlow works out-of-the-box without tuning across multiple domains, including large-scale autonomous driving scenes and dynamic tabletop settings. Remarkably, EulerFlow produces high quality flow estimates on small, fast moving objects like birds and tennis balls, and exhibits emergent 3D point tracking behavior by solving its estimated PDE over long-time horizons. On the Argoverse 2 2024 Scene Flow Challenge, EulerFlow outperforms all prior art, surpassing the next-best unsupervised method by more than 2.5×2.5$\times$2.5 ×, and even exceeding the next-best supervised method by over 10%.
@@
Figure 1: EulerFlow is able to capture the motion of small, fast moving objects with few lidar points, such a bird flying in front of an autonomous vehicle (Figure 1(a)). EulerFlow’s flexibility allows it to estimate scene flow for fast-moving table top objects without additional hyperparameter tuning (Figure 1(b)). EulerFlow’s PDE estimate exhibits emergent 3D point tracking behavior without explicit long-horizon supervision (Figure 1(c)). Note that point clouds are shown in color for visualization purposes only; RGB is not used during optimization.
Interactive scene visualizations at vedder.io/eulerflow
@@ 1 Introduction
Scene flow estimation is the task of describing motion with per-point 3D motion vectors between temporally successive point clouds (Dewan et al., 2016; Liu et al., 2019; Erçelik et al., 2022; Jund et al., 2021; Zhang et al., 2024b; Vedder et al., 2024; Khatri et al., 2024). Such per-point motion estimates are critical for autonomy in diverse environments, e.g. maneuvering around open-world objects like flying debris (Peri et al., 2022a) or folding deformable cloth (Weng et al., 2022). Importantly, scene flow estimation requires not only an understanding of object geometry, but also its motion. However, scene flow methods broadly do not work on smaller objects (Khatri et al., 2024). For example, in the autonomous vehicles domain, Khatri et al. highlight that even supervised methods struggle to describe the majority of pedestrian motion, with unsupervised methods failing dramatically. Scene flow promises to be a powerful primitive for understanding the dynamic world, but such failures explain why it has limited adoption in downstream applications like tracking (Zhai et al., 2020) or open-world object extraction (Najibi et al., 2022).
@@
Scene Flow via PDE. In Figure 2, visual assessment of scene flow quality is much easier in an accumulated global frame; while incomplete due to an implicit time axis, these accumulated flow vectors allow viewers to imagine how positions in 3D space evolve over many timesteps, and compare that to predicted flows. This imagination of scene flow as continuous motion over large time intervals motivates us to model scene flow as a Partial Differential Equation (PDE) that describes the scene’s instantaneous motion across continuous position and time; this is the same formalism used to represent fluid flow or heat flow in classical mechanics. Scene flow estimation then becomes the task of estimating this PDE. We can straightfowardly represent this PDE estimate with a neural prior (Li et al., 2021b) and optimize it against scene flow surrogate objectives, both over single frame pairs and extended across arbitrary time intervals, unlocking new optimization objectives that produce better quality estimates. We formalize this in Section 3 and propose the Scene Flow via PDE framework.
EulerFlow. We instantiate Scene Flow via PDE with standard point cloud distance objectives like Chamfer Distance to create EulerFlow. Notably, EulerFlow outperforms all prior art, supervised or unsupervised, on the Argoverse 2 2024 Scene Flow Challenge and Waymo Open Scene Flow benchmark. It outperforms prior unsupervised methods by a large margin (>2.5×>2.5$\times$> 2.5 × mean dynamic error reduction), and is able to capture small, fast moving objects, including those outside of labeled taxonomies (e.g. the flying bird in Figure 1(a)). Due to its simplicity, EulerFlow is able to provide high quality scene flow out-of-the-box on real-world data for other important domains such as dynamic tabletop settings (Figure 1(b)) without domain-specific tuning. Finally, we show that simple PDE solving techniques like Euler integration can be used to extract 3D point tracking behaviors (Figure 1(c)), which serves as both an exciting emergent behavior as well as a mechanism for visualizing and interpreting the quality of the continuous PDE estimate.
@@
    We propose Scene Flow via PDE (SFvPDE), a reframing of scene flow estimation as the task of fitting a PDE that describes the change of continuous positions over continuous time, unlocking a new class of surrogate objectives that enable better scene flow estimates.
@@
    We instantiate SFvPDE with EulerFlow, a flexible unsupervised scene flow method that achieves state-of-the-art performance on the Argoverse 2 2024 Scene Flow Challenge, beating all prior supervised and unsupervised methods.
@@
    We study EulerFlow and show its strong performance is derived from its ability to optimize its PDE estimate against the full sequence of observations over arbitrary time horizons.
@@ 2 Background and Related Work
Evaluation. Dewan et al. formalized scene flow for point clouds as the task of estimating motion between point cloud $P_{t}{}$ at time $t$ and point cloud $P_{t+1}{}$ at $t+1$ by estimating the true flow $F^{*}_{t,t+1}$, i.e. true residual vectors for every point in $P_{t}{}$ that describe its movement to its associated position at $t+1$. Error is computed by measuring the per-point endpoint distance between estimated and ground truth vectors. Historically, these errors are reported with a per-point average (Average EPE); however, as Chodosh et al. show, Average EPE is dominated by background points, preventing meaningful measurement of non-ego object motion descriptions. Khatri et al. address this shortcoming with Bucket Normalized EPE, which reports per-class performance normalized by speed, allowing for direct comparisons across classes with very different average speeds (e.g. pedestrians and cars). Bucket Normalized EPE is the basis for the Argoverse 2024 Scene Flow Challenge [^1] https://www.argoverse.org/sceneflow, where methods are ranked by the mean error of their motion descriptions (mean Dynamic Normalized EPE).
Input / Output Formulation. Dewan et al.’s choice to formulate scene flow using only two input frames is arbitrary; it’s the minimal information needed to extract rigid motion, but there are not real-world problems constrained to only have access to two frames. Indeed, using five or ten frames of past observations is standard practice in the 3D detection literature (Zhu et al., 2019; Vedder & Eaton, 2022; Peri et al., 2022b; 2023; Nalty et al., 2022), and multi-frame formulations have started to appear in the scene flow literature: Liu et al. (2024) and Flow4D (Kim et al., 2024) use three ($P_{t-1},P_{t},P_{t+1}$) and five input frames ($P_{t-3},\ldots,P_{t+1}$) respectively to predict $\hat{F}_{t,t+1}$. As we discuss in Section 3, rather than just using more observations to estimate flow for a single frame pair, we formulate scene flow as a joint estimation problem: given the full observation sequence $\left(P_{0},\ldots,P_{N}\right)$, we estimate all flows $\hat{F}_{0,1},\ldots,\hat{F}_{N-1,N}$ between all adjacent observations.
@@
Neural Scene Flow Prior. Li et al. (2021b) propose Neural Scene Flow Prior (NSFP), a widely adopted unsupervised scene flow approach. NSFP uses the inductive bias of the smooth, restricted learnable function class of two ReLU MLP coordinate networks (8 hidden layers of 128 neurons; one to estimate forward flow $\hat{F}^{+}$ and to estimate reverse flow $\hat{F}^{-}$) to minimize
@@
Deformation in Reconstruction. Nerfies (Park et al., 2021) and DynamicFusion (Newcombe et al., 2015) estimate a deformation field to warp a canonical frame to explain the observed frame. While capable of describing small motions, these methods require a canonical frame that contains all of the relevant geometry to deform; however, in large, highly dynamic scenes like autonomous driving, there is often no frame that contains all moving objects. By comparison, Scene Flow via PDE does not assume the existence of a canonical frame, instead only describing how the scene changes.
@@
Figure 3: Overview of our Scene Flow via PDE framework, which estimates a position PDE across the entire observation sequence by optimizing against multi-frame objectives. This PDE estimate is represented with a neural prior (Li et al., 2021b), providing a flexible, general representation for describing position-time motion.
3 Scene Flow via PDE
@@
Rather than consuming an arbitrary history of observations to estimate the true flow for a single adjacent observation pair, we pose the problem of jointly estimating the true flow $F^{*}_{t,t+1}$ for all adjacent point clouds $P_{t},P_{t+1}$ in the entire sequence $\left(P_{0},\ldots,P_{N}\right)$. To do this, rather than describing scene flow as positional change over a fixed interval ($F^{*}_{t,t+1}$ are residual vectors over the interval $t$ to $t+1$) as we did in Section 2, we can instead express these changes as a differential equation that describes positional change over continuous time.
Given some scene, let $L^{*}{}\left(\mathbf{x};t\right)$ be the continuous function that describes the true position of coordinate $\mathbf{x}\in\mathbb{R}^{3}$ from a canonical frame at some arbitrary time $t\in\mathbb{R}$. The function describing how locations change over time would be $L^{*}$’s partial derivative, the Partial Differential Equation $\frac{\partial L^{*}}{\partial t}$. This formulation allows us to recover flow vectors over any arbitrary continuous time interval; for example, solving the PDE over the time interval $t$ to $t+1$ with $\vec{0}$ initial conditions and querying that function with the points of $P_{t}$ will recover $F^{*}_{t,t+1}$. Additionally, it also avoids the important limitation of requiring a canonical frame; the canonical frame we describe defining $L^{*}$ disappears during differentiation, as $\frac{\partial L^{*}}{\partial t}$ merely describes its change.
@@
While we do not have access to the true $L^{*}$ or its PDE $\frac{\partial L^{*}}{\partial t}$, we can estimate it by optimizing a single neural prior $\theta$ against surrogate objectives. This framing, which we formalize into the Scene Flow via PDE framework (SFvPDE; Figure 3), allows the representation of $\theta$ to benefit from constructive interference between objectives, as well as formulate objectives over arbitrarily long time horizons, enabling an optimization driven estimate of $\frac{\partial L^{*}}{\partial t}$ (Equation 2).
$\frac{\partial L^{*}}{\partial t}\simeq\textup{SFvPDE{}}(\theta,t)$ (2)
In addition to not requiring a canonical frame definition, this formulation also benefits from the use of ODE solvers to extract residual vectors rather than regressing them directly. For example, in the deformation fields described in Section 2, the deformation estimator must directly output highly non-linear warps of the canonical frame across arbitrary time. In contrast, SFvPDE only requires that methods estimate local derivatives, with significant non-linear expressivity provided by the ODE solver’s numerical integration.
@@ 4 EulerFlow
We straightforwardly instantiate Scene Flow via PDE with EulerFlow, a point cloud only scene flow method [^2] Visualizations shown in color for better viewing. EulerFlow can also use monodepth estimates (Supp. A.2) . EulerFlow is far simpler that prior art, requiring only a single optimization loop over a single neural prior. Our neural prior $\theta$ is a straightforward extension to NSFP’s coordinate network prior [^3] Hyperparameters (e.g. filter width of 128) of NSFP’s prior are kept fixed, except for depth (Section 5.2.3).; however, instead of taking a 3D space vector (positions $X,Y,Z\in\mathbb{R}$) as input, we encode a 5D space-time-direction vector: positions $X,Y,Z,\in\mathbb{R}$, sequence normalized time $t\in[-1,1]$ (i.e. the point cloud time scaled to this range), and direction $d\in\{\texttt{BWD}=-1,\texttt{FWD}=1\}$. This simple encoding scheme enables description of arbitrary regions of the PDE, allowing for the PDE to be queried at frequencies different from the sensor frame rate. Euler integration enables simple implementation of multi-step forward, backward, and cyclic consistency losses without extra bells and whistles. For efficiency, we use Euler integration with $\Delta{}t$ set as the time between observations for our PDE solver, enabling support for arbitrary sensor frame rates. We optimize $\theta{}$ as formulated in Equation 3, featuring bi-directional three step Euler integration and one step cycle consistency, over 1000 epochs using Adam optimizer with a learning rate $8e^{-5}$ with five frame minibatches and NSFP’s early stopping schedule. Formally, for $P_{t}$ we optimize
$\arg\min_{\theta}\sum{\begin{array}[]{l}\forall k\in\{1,2,3\}:\textup{TruncatedChamfer{}}(\textup{Euler}_{\theta}\left(P_{t},\texttt{FWD},k\right),P_{t+k})\\ \forall k\in\{1,2,3\}:\textup{TruncatedChamfer{}}(\textup{Euler}_{\theta}\left(P_{t},\texttt{BWD},k\right),P_{t-k})\\ \left\lVert\textup{Euler}_{\theta}\left(\textup{Euler}_{\theta}\left(P_{t},\texttt{FWD},1\right),\texttt{BWD},1\right)-P_{t}\right\rVert_{2}\times 0.01\end{array}}\enspace.$ (3)
Like with NSFP, TruncatedChamfer is defined as the standard $L_{2}$ Chamfer distance with per-point distances below 2 meters. Additionally, we introduce $\textup{Euler}_{\theta}\left(P_{t},d,k\right)$, which is $k$ step Euler integration from $t$ to $t\pm k$ depending on the direction $d$, with the given point cloud $P_{t}$ as the initial conditions. As we show in Section 5, EulerFlow’s simple PDE estimation formulation produces high quality flow and solving this PDE over arbitrary time spans unlocks emergent point tracking behavior.

### v2

# Neural Eulerian Scene Flow Fields

@@ Abstract
We reframe scene flow as the task of estimating a continuous space-time ordinary differential equation (ODE) that describes motion for an entire observation sequence, represented with a neural prior. Our method, EulerFlow, optimizes this neural prior estimate against several multi-observation reconstruction objectives, enabling high quality scene flow estimation via self-supervision on real-world data. EulerFlow works out-of-the-box without tuning across multiple domains, including large-scale autonomous driving scenes and dynamic tabletop settings. Remarkably, EulerFlow produces high quality flow estimates on small, fast moving objects like birds and tennis balls, and exhibits emergent 3D point tracking behavior by solving its estimated ODE over long-time horizons. On the Argoverse 2 2024 Scene Flow Challenge, EulerFlow outperforms all prior art, surpassing the next-best unsupervised method by more than 2.5×2.5$\times$2.5 ×, and even exceeding the next-best supervised method by over 10%. See vedder.io/eulerflow for interactive visuals.
@@
Figure 1: EulerFlow is able to capture the motion of small, fast moving objects with few lidar points, such a bird flying in front of an autonomous vehicle (Figure 1(a)). EulerFlow’s flexibility allows it to estimate scene flow for fast-moving table top objects without additional hyperparameter tuning (Figure 1(b)). EulerFlow’s ODE estimate exhibits emergent 3D point tracking behavior without explicit long-horizon supervision (Figure 1(c)). Note that point clouds are shown in color for visualization purposes only; RGB is not used during optimization.
@@ 1 Introduction
Scene flow estimation is the task of describing motion with per-point 3D motion vectors between temporally successive point clouds (Dewan et al., 2016; Liu et al., 2019; Erçelik et al., 2022; Jund et al., 2021; Zhang et al., 2024b; Vedder et al., 2024; Khatri et al., 2024). Such per-point motion estimates are critical for autonomy in diverse environments, e.g., maneuvering around open-world objects like debris (Peri et al., 2022a) or folding deformable cloth (Weng et al., 2022). Importantly, scene flow estimation requires not only an understanding of object geometry, but also its motion. However, scene flow methods broadly do not work on smaller objects (Khatri et al., 2024). For example, in the autonomous vehicles domain, Khatri et al. highlight that even supervised methods struggle to describe the majority of pedestrian motion, with unsupervised methods failing dramatically. Scene flow promises to be a powerful primitive for understanding the dynamic world, but such failures explain why it has limited adoption in downstream applications like tracking (Zhai et al., 2020) or open-world object extraction (Najibi et al., 2022).
@@
Scene Flow via ODE. In Figure 2, visual assessment of scene flow quality is much easier in an accumulated global frame; while incomplete due to an implicit time axis, these accumulated flow vectors allow viewers to imagine how positions in 3D space evolve over many timesteps, and compare that to predicted flows. This imagination of scene flow as continuous motion over large time intervals motivates us to model scene flow as an ordinary differential equation (ODE) that describes the scene’s instantaneous motion across continuous position and time. Scene flow estimation then becomes the task of estimating this ODE. We can straightfowardly represent this ODE estimate with a neural prior (Li et al., 2021b) and optimize it against scene flow surrogate objectives, both over single frame pairs and extended across arbitrary time intervals to produce better quality estimates. We formalize this in Section 3 and propose the Scene Flow via ODE framework.
EulerFlow. We instantiate Scene Flow via ODE with standard point cloud distance objectives like Chamfer Distance to create EulerFlow. Notably, EulerFlow outperforms all prior art, supervised or unsupervised, on the Argoverse 2 2024 Scene Flow Challenge and Waymo Open Scene Flow benchmark. It outperforms prior unsupervised methods by a large margin (>2.5×>2.5$\times$> 2.5 × mean dynamic error reduction), and is able to capture small, fast moving objects, including those outside of labeled taxonomies (e.g. the flying bird in Figure 1(a)). Due to its simplicity, EulerFlow is able to provide high quality scene flow out-of-the-box on real-world data for other important domains such as dynamic tabletop settings (Figure 1(b)) without domain-specific tuning. Finally, we show that simple ODE solving techniques like Euler integration can be used to extract 3D point tracks (Figure 1(c)), which serves as both an exciting emergent behavior as well as a mechanism for visualizing and interpreting the quality of the continuous ODE estimate.
@@
    We propose Scene Flow via ODE (SFvODE), a reframing of scene flow estimation as the task of fitting an ODE that describes the change of continuous positions over continuous time, unlocking a new class of surrogate objectives that enable better scene flow estimates.
@@
    We instantiate SFvODE with EulerFlow, a flexible unsupervised scene flow method that achieves state-of-the-art performance on the Argoverse 2 2024 Scene Flow Challenge, beating all prior supervised and unsupervised methods.
@@
    We study EulerFlow and show its strong performance is derived from its ability to optimize its ODE estimate against the full sequence of observations over arbitrary time horizons.
@@ 2 Background and Related Work
Evaluation. Dewan et al. formalized scene flow for point clouds as the task of estimating motion between point cloud $P_{t}{}$ at time $t$ and point cloud $P_{t+1}{}$ at $t+1$ by estimating the true flow $\mathcal{F}_{t,t+1}$, i.e. true residual vectors for every point in $P_{t}{}$ that describe its movement to its associated position at $t+1$. Error is computed by measuring the per-point endpoint distance between estimated and ground truth vectors. Historically, these errors are reported with a per-point average (Average EPE); however, as Chodosh et al. show, Average EPE is dominated by background points, preventing meaningful measurement of non-ego object motion descriptions. Khatri et al. address this shortcoming with Bucket Normalized EPE, which reports per-class performance normalized by speed, allowing for direct comparisons across classes with very different average speeds (e.g. pedestrians and cars). Bucket Normalized EPE is the basis for the Argoverse 2024 Scene Flow Challenge [^1] https://www.argoverse.org/sceneflow, where methods are ranked by the mean error of their motion descriptions (mean Dynamic Normalized EPE).
Input / Output Formulation. Dewan et al.’s choice to formulate scene flow using only two input frames is arbitrary; it’s the minimal information needed to extract rigid motion, but there are not real-world problems constrained to only have access to two frames. Indeed, using five or ten frames of past observations is standard practice in the 3D detection literature (Zhu et al., 2019; Vedder & Eaton, 2022; Peri et al., 2022b; 2023; Nalty et al., 2022), and multi-frame formulations have started to appear in the scene flow literature: Liu et al. (2024) and Flow4D (Kim et al., 2024) use three ($P_{t-1},P_{t},P_{t+1}$) and five input frames ($P_{t-3},\ldots,P_{t+1}$) respectively to predict ^$\hat{\mathcal{F}}_{t,t+1}$. As we discuss in Section 3, rather than just using more observations to estimate flow for a single frame pair, we formulate scene flow as a joint estimation problem: given the full observation sequence $\left(P_{0},\ldots,P_{N}\right)$, we estimate all flows ^$\hat{\mathcal{F}}_{0,1},\ldots,\hat{\mathcal{F}}_{N-1,N}$ between all adjacent observations.
@@
Neural Scene Flow Prior. Li et al. (2021b) propose Neural Scene Flow Prior (NSFP), a widely adopted unsupervised scene flow approach. NSFP uses the inductive bias of the smooth, restricted learnable function class of two ReLU MLP coordinate networks (8 hidden layers of 128 neurons); $\theta$ to estimate forward flow and $\theta^{\prime}$ to estimate backwards flow, minimizing
@@
Deformation in Reconstruction. Nerfies (Park et al., 2021) and DynamicFusion (Newcombe et al., 2015) estimate a deformation field to warp a canonical frame to explain the observed frame. While capable of describing small motions, these methods require a canonical frame that contains all of the relevant geometry to deform; however, in large, highly dynamic scenes like autonomous driving, there is often no frame that contains all moving objects. By comparison, Scene Flow via ODE does not assume the existence of a canonical frame, instead only describing how the scene changes.
@@
Figure 3: Overview of our Scene Flow via ODE framework, which estimates an ODE across the entire observation sequence by optimizing against multi-frame objectives. This ODE estimate is represented with a neural prior (Li et al., 2021b), providing a flexible, general representation for describing position-time motion.
3 Scene Flow via ODE
@@
Prior art consumes multiple frames $\left(P_{t-N},\ldots,P_{t+1}\right)$ as input, but these methods are ultimately only tasked with estimating flow vectors between $P_{t}$ and $P_{t+1}$. We instead pose the problem of estimating a time-conditioned flow field that describes motion for all adjacent point clouds $P_{t},P_{t+1}$ in the entire sequence $\left(P_{0},\ldots,P_{N}\right)$. To do this, rather than describing scene flow as positional change over a fixed interval ($\mathcal{F}_{t,t+1}$ are residual vectors over the interval $t$ to $t+1$) as we did in Section 2, we can instead express these changes as a differential equation that describes positional change over continuous time.
[…]
Figure 4: Comparison of Eulerian and Lagrangian descriptions of 2D flow. An Eulerian view characterizes a flow field via instantaneous velocities at many different points, while a Lagrangian view characterizes a flow field via trajectories of many different particles across time. Both approaches are valid ways of describing an underlying flow field, and with sufficient characterization one view can be readily converted to another, but the Lagrangian view relies on a definition of the definition of consistent canonical frame.
Formally, given a scene, let $L(x_{0},y_{0},z_{0},t)$ be the Lagrangian view of the scene’s true flow field, i.e. a continuous function that, based on a canonical frame at time 00, describes the true position of the canonical frame particle $x_{0},y_{0},z_{0}$ at some other time $t$. As we discuss in Section 2, this Lagrangian view is common in the the deformable reconstruction literature, and the requirement for a canonical frame definition means these approaches struggle to describe scenes where there is no frame that contains all moving objects.
To break this canonical frame dependence, we choose to take an Eulerian view of the flow field, i.e. $F=\frac{dL}{dt}$, which describes the velocity of a query point at some arbitrary time. As we show in our derivation in Appendix D, this formulation does not require point correspondences in some other canonical frame when estimating a point’s trajectory from $t$ to $t^{\prime}$; instead, we can simply set the initial conditions of the ODE at $t$ to $x_{t},y_{t},z_{t}$ and utilize an off-the-shelf ODE solver (e.g. Euler integration) to extract flow from $t$ to $t^{\prime}$, expressed as $E(x_{t},y_{t},z_{t},t,t^{\prime})$.
@@
We do not know the true flow field $F$ when estimating scene flow; however, we can represent $F$ with a neural prior $\theta$ ($F\approx\theta$), and optimize $\theta$ against surrogate objectives. This framing, which we formalize into the Scene Flow via ODE framework (SFvODE; Figure 3), allows $\theta$ to benefit from constructive interference between objectives, as well as enables us to formulate objectives over arbitrarily long time horizons, unlocking high quality estimates.
@@ 4 EulerFlow
Scene Flow via ODE proposes a framework where the neural prior $\theta$ represents an estimate of the Eulerian flow field $F$ (i.e. $F\approx\theta$); however, it does not prescribe the optimization objectives for $\theta$. Thus, we instantiate Scene Flow via ODE with EulerFlow, a point cloud only scene flow method [^2] Visualizations shown in color for better viewing. EulerFlow can also use monodepth estimates (Appendix A.2) with reconstruction and cyclic consistency objectives across the entire sequence of observations.
As we show in Equation 17 (Appendix D.4), we can use $\theta$’s Eulerian flow field estimate to extract an estimated point trajectory from $x_{t},y_{t},z_{t}$ at $t$ to some future location at time $t^{\prime}$ via Euler integration over $\theta$ without requiring a canonical frame definition, i.e. $E_{\theta}(x_{t},y_{t},z_{t},t,t^{\prime})$. By extracting point trajectories for every point $p$ in $P_{t}$ using $E_{\theta}$, we can not only construct a two-frame scene flow estimate of $\mathcal{F}_{t,t+1}$, but also estimate flow to arbitrary future or prior timesteps (e.g. $\mathcal{F}_{t,t+2}$ or $\mathcal{F}_{t,t-1}$). This allows us to optimize over multi-frame reconstruction objectives: we can now pose reconstruction surrogate objectives between any two point clouds in our observation sequence, not just adjacent point clouds $P_{t}$ and $P_{t+1}$. Similarly, we can straightforwardly pose cyclic consistency objectives by composing $\mathcal{F}_{t,t+1}$ and $\mathcal{F}_{t+1,t}$. Formally, for $P_{t}$’s $\mathcal{F}_{t,t+k}$ (for any $k\in\mathbb{Z}$), we define
$\textup{Euler}_{\theta}\left(P_{t},k\right)=P_{t}+\mathcal{F}_{t,t+k}=\forall p\in P_{t}:E_{\theta}(p_{xt},p_{yt},p_{zt},t,t+k)\enspace,$ (2)
enabling us to pose $\theta$’s optimization objective $\forall P_{t}\in\left(P_{0},\ldots,P_{N}\right)$ across the window of size $W$
$\arg\min_{\theta}\sum{\begin{array}[]{l}\forall k\in\{-W,\ldots,W\}\setminus\{0\}:\textup{TruncatedChamfer{}}(\textup{Euler}_{\theta}\left(P_{t},k\right),P_{t+k})\\ \alpha\left\lVert\textup{Euler}_{\theta}\left(\textup{Euler}_{\theta}\left(P_{t},1\right),-1\right)-P_{t}\right\rVert_{2}\end{array}}\enspace$ (3)
In practice, we set W to 3 and $\alpha$ to $0.01$. We provide additional implementation details in Appendix C. In order to optimize $\theta$, our estimate of the Eulerian flow field $F$, we perform Euler integration to extract point cloud flow estimates as part of reconstruction losses. Notably, EulerFlow only requires a single optimization loop over a single neural prior $\theta$ compared to NSFP’s two priors $\theta$ and $\theta^{\prime}$. Our neural prior $\theta$ is a straightforward extension to NSFP’s coordinate network prior. Like with NSFP, TruncatedChamfer is defined as the standard $L_{2}$ Chamfer distance with per-point distances below 2 meters. As we show in Section 5, EulerFlow’s simple ODE estimation formulation across multiple observations produces high quality flow, and solving this ODE over arbitrary time spans unlocks emergent point tracking behavior.